# LightRetriever: A LLM-based Text Retrieval Architecture with Extremely Faster Query Inference

### v4

@@ 4 Ablations
Our core design adopts an asymmetric architecture, with a lightweight query MLP Embedding and a full-sized document model for inference. Specifically, we utilize a full-sized model on both sides at training time, then cache all query token embeddings in one Embedding Lookup for fast query-side inference. In ablation studies below (A1-A3), we are interested in other potential symmetric/asymmetric architectures, which are described in the following Table 4.
Table 3: Symmetry Ablation settings (A1-A3). 1L-Trans denotes a single-layer Transformer.
@@
Ours Full Full MLP Full
A1 Full Full MLP MLP
A2 MLP Full MLP Full
A3 1L-Trans Full 1L-Trans Full
Table 4: Ablations (A1-3) results (nDCG@10) of LightRetriever.
@@
To test the necessity of the above asymmetry, we evaluate a symmetric lightweight setup where both sides use simplified dense encoders. As shown in Table 4, this leads to severe performance drops on both BeIR (–13.8) and CMTEB-R (–18.6), confirming that full document representations are crucial for maintaining retrieval effectiveness. This symmetric setting is not functional for sparse retrieval due to the absence of learnable parameters.
The best symmetric lightweight retriever is Static Embedding (one MLP) + BM25 (term-based) in Table 2, with nDCG@10 of 44.7 on BeIR and 52.1 on CMTEB-R. However, due to the lack of deep modeling on documents, such a combination is still outperformed.
We also tested a reversed asymmetry with deep encoding for queries and lightweight documents. However, we observed training instability and degraded performance, likely caused by embedding collapse.
A2. Can we use only a query MLP for training?
To confirm that a deep query encoder is still needed during training, we replaced the Transformer-based query encoder with a simple MLP that reuses only the LLM’s embedding layer. This setup significantly degrades performance (–11.2 on BeIR, –17.4 on CMTEB-R), indicating that the full-sized modeling for query side training remains essential during training to learn effective token-level representations.
A3. Is adding one Transformers layer better on the query side?
We also evaluated using only the first Transformer layer on the query side. While this reduces model size during training, it still underperforms the LightRetriever (–4.3 on BeIR, –8.6 on CMTEB-R) and introduces inference-time overhead, making it incompatible with the design goals of LightRetriever. This again proves the need for full-sized query modeling during the training phase.
Figure 5: Ablations of retrieval performances (nDCG@10) of LightRetriever on BeIR and CMTEB-R benchmarks.
A4. Breakdown Ablation of Retrieval Performance
Figure 5 reports LightRetriever’s performance under dense, sparse, and hybrid configurations. Dense retrieval consistently outperforms sparse across all backbones (e.g., LLaMA3, Qwen2.5) and model sizes (1B–8B).
While sparse retrieval alone is less effective, the hybrid approach recovers most of the performance gap with minimal overhead. For instance, LLaMA3-1B achieves 52.0 and 60.8 nDCG@10 on BeIR and CMTEB-R in the hybrid setting, with only 1.1 and 3.0 points below the full model, but with a much smaller query encoder. These results confirm the complementarity of both signals and the effectiveness of LightRetriever in balancing efficiency and accuracy.
@@ 5 Conclusion
Existing LLM-based retrievers use symmetric dual-encoders to model both queries and documents. While documents could be pre-encoded and indexed offline, queries arrive in real-time and need online encoding. The deployment of large LLMs results in inefficient online services. In this paper, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture capable of extremely lightweight query encoding. Our approach achieves over a 1000x speedup for query inference on an H800 GPU, even a 20x speedup without GPU acceleration. Experiments show great robustness and generalization abilities across different foundation LLMs and retrieval tasks.
Appendix A Technical Appendices and Supplementary Material
@@ A.1 Limitations
Our work focuses on online efficiency optimizations of recent LLM-based architecture, whose giant parameters pose a greater need for such optimizations. Other readers may be interested in the application of our methods on non-LLM architectures, e.g., BERT. While such a combination is possible, due to limited time constraints and lesser needs for non-LLM, we leave this to future work.
@@ A.2.1 Text retrieval
Text retrieval builds on top of the vector space model [63], which encodes query and document representation vectors, and performs relevance search based on vector similarities.
@@
Dense retrieval [30] trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) [52] algorithm. In recent years, dense retrievers, such as BGE [11] and E5 [71], have gained popularity for strong retrieval abilities across different tasks [51, 66], languages [84], and context granulaties [75, 53], owing to diverse training data [60], retrieval-customized pre-training [22, 76, 47], improved negative-mining [79], and so on. Recent SOTA retrievers [72, 5, 39] start to utilize LLMs [27, 68] as backbone encoders. With larger parameters and pre-training data, LLM-based retrievers enable significantly powerful retrieval capacities and domain adaptation abilities. However, LLMs also incur heavy online workloads when serving as query encoders. Improving the serving efficiency of LLM-based retrievers is still left for exploration.
@@
Traditional statistical sparse retrieval algorithms (e.g., TF-IDF [64], and BM25 [61]) do not involve learnable language models. Instead, they directly analyze lexical frequencies (e.g., TF) and importance (e.g., IDF) in a large vocabulary space. These lexicals are sparsely distributed and require building inverted indexes [46] for feasible similarity metric calculation, such as Apache Lucene [6] and Tantivy [59]. In recent years, sparse retrieval has started to work with pre-trained LMs (PLMs). BGE [11], SPLADE [21], SparTerm [3], and SPARTA [85] use customized XLM-RoBERTa [14] or BERT [19] as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors. Because of sparsity, these works also search with inverted indexes. Recent works [86] also try to incorporate LLMs with sparse retrieval. However, the online efficiency issue still exists.
@@
Hybrid retrieval is a technique that combines retrieval scores [73] or ranks [15] from multiple retrieval systems, such as dense [30] / sparse [61, 21] retrieval, and single [30] / multiple [32, 11] vector retrieval. It enables better retrieval abilities by interpolating results from multiple sources. Linear interpolation [73] of dense and sparse scores is used in our work.
@@ A.2.2 Improving inference efficiency
Previous works have tried to distill small query encoders from large encoders for better query inference efficiency. KALE [9] prunes the BERT-based query encoder and distills it during post-training. However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ [37] decreases by 25-30%. [74] proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models. Its experiments show that a 2-layer BERT query encoder retains 92.5% of the full-sized performances via distillation and proper student initialization. USTAD [33] also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students. Its experiments show that the distilled BERT-mini query encoder achieves 95-97% of the teacher’s performance because of better distillation.
However, these studies have primarily focused on small language models (SLMs), such as BERT [19], and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical. Furthermore, these existing methods typically retain Transformer layers in the query encoders, which limits their efficiency gains. In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO [53], thus unable to compete with LLM-based retrievers.
@@
Previous works have also been dedicated to reducing index sizes for smaller disk space usage. Matryoshka representation learning (MRL) [36] supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors. It aligns and trains the corresponding multi-dimensional vectors in the end-to-end contrastive learning. For example, if $k\in\{64,128,256\}$, query-passage dense vectors with $\{64,128,256\}$-dims could be obtained together after one encoder forward operation. For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals [21]. Thus, it naturally supports reducing the index sizes by retaining Top-k terms/lexicals. To explore the potential of controlling index sizes, our work also explores shrinking index sizes with dense MRL and sparse Top-k in the Section A.3, which also shows strong generalization capacities.
@@
A.3 Additional results about controls of vector dimension and sparsity
Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors. For example, indexing 8.8M MS-MARCO passages [53] with Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval. Thus, it’s essential to explore the possibility of controlling index sizes by shrinking vectors. 1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) [36] with LightRetriever to support flexible controls of dense vector dimensions. Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors. MRL trains a collection of Top-k dimensions together, enabling flexible choosing of a proper k. 2) For sparse retrieval, the sparse vector directly learns the term impact (i.e., lexical frequency) within a large sparse dimension. Thus, LightRetriever supports controlling sparsity by directly taking the Top-k terms, where the larger impact terms matter more.
We conduct experiments of controlling dense vector dimension and sparse vector sparsity with Top-k of $\{128,256,512,1024,2048,4096\}$ on Llama-3.1-8b. As presented in Figure 6, the performances of LightRetriever decrease marginally when shrinking Top-k to 4x ($4090\rightarrow 1024$). The lowest Top-k setting ($128$) has 7.9% and 8.8% performance gaps on BeIR and CMTEB-R benchmarks compared to the highest Top-k ($4096$), but its disk usages with MS-MARCO are only 4.22 GB for dense trieval and 3.03 GB for sparse retrieval. This shows that our LightRetriever is capable of flexible index controls via simple vector shrinking techniques, with limited effect on retrieval performances.
A.4 Additional ablation about the effect of auxiliary KL loss

### v5

@@ 4 Ablations
Our core design adopts an asymmetric architecture, with a lightweight query Embedding Bag (Emb) and a full-sized document model for inference. Specifically, we utilize a full-sized model on both sides at training time, then cache all query token embeddings in one Embedding Lookup for fast query-side inference. In ablation studies below (A1-A2), we are interested in other potential symmetric/asymmetric architectures, which are described in the following Table 5.
Table 4: Symmetry Ablation settings (A1-A2). Emb denotes Embedding Bag. 1L-Trans denotes a single-layer Transformer.
@@
Ours Full Full Emb Full
A1 Full Full Emb Emb
A2 Emb Full Emb Full
Table 5: Ablations (A1-2) results (nDCG@10) of LightRetriever.
@@
To test the necessity of above asymmetry, we evaluate a symmetric lightweight setup where both sides use simplified dense encoders. As shown in Table 5, this leads to severe performance drops on both BeIR (–13.8) and CMTEB-R (–18.6), confirming that full document representations are crucial for maintaining retrieval effectiveness. This symmetric setting is not functional for sparse retrieval due to the absence of learnable parameters.
The best symmetric lightweight retriever is Static Embedding (one Embedding Bag) + BM25 (term-based) in Table 2, with nDCG@10 of 44.7 on BeIR and 52.1 on CMTEB-R. However, due to the lack of deep modeling on documents, such a combination is still outperformed. We also tested a reversed asymmetry with deep query encoding and lightweight documents. However, we observed training instability and degraded performance, likely caused by embedding collapse.
A2. Can we use only a query Embedding Bag for training?
To confirm that a deep query encoder is still needed during training, we replaced the Transformer-based query encoder with a simple Embedding Bag that reuses only the LLM’s embedding layer. This setup significantly degrades performance (–11.2 on BeIR, –17.4 on CMTEB-R), indicating that although deep token interaction could be limited with a durable cost during query inferencing, the full-sized modeling for query side training remains essential to learn effective representations.
Figure 5: Ablations of retrieval performances (nDCG@10) of LightRetriever on BeIR and CMTEB-R benchmarks. For detailed results, please refer to Table 14, 15 and 16 in Appendix.
A3. Breakdown Ablation of Retrieval Performance
Figure 5 reports LightRetriever’s performance under dense, sparse, and hybrid configurations. Dense retrieval consistently outperforms sparse across all backbones (e.g., LLaMA3, Qwen2.5) and model sizes (1B–8B). While sparse retrieval alone is less effective, the hybrid approach recovers most of the performance gap with minimal overhead. For instance, LLaMA3-1B achieves 52.0 and 60.8 nDCG@10 on BeIR and CMTEB-R in the hybrid setting, with only 1.1 and 3.0 points below the full model, but with a much smaller query encoder. These results confirm the complementarity of both signals and the effectiveness of LightRetriever in balancing efficiency and accuracy.
@@ 5 Conclusion
Existing LLM-based retrievers use symmetric dual-encoders to model both queries and documents. While documents could be pre-encoded and indexed offline, queries arrive in real-time and need online encoding. The deployment of LLMs results in inefficient online services. In this paper, we propose LightRetriever, a novel LLM-based hybrid retrieval architecture capable of extremely lightweight query encoding. Our approach achieves over a 1000x speedup for query inference and 10x increase in end-to-end retrieval throughput on an A800 GPU. Experiments show great robustness and generalization abilities across different foundation LLMs and retrieval tasks.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. U24A20335).
Appendix A Appendix
@@ A.1 Limitations
Our work focuses on online efficiency optimizations of recent LLM-based architectures, whose giant parameters pose a greater need for such optimizations. Other readers may be interested in the application of our methods on non-LLM architectures, e.g., BERT. While such a combination is possible, due to limited time constraints and lesser needs for non-LLM, we leave this to future work. Moreover, as stated in Section §3.2, our query-lightweight design is applicable to normal relevance-based retrieval tasks. Tasks that require complex query understanding, such as reasoning-intensive retrieval, are not suitable for our work.
@@ A.2.1 Text retrieval
Text retrieval builds on top of the vector space model (Salton et al., 1975), which encodes query and document representation vectors, and performs relevance search based on vector similarities.
@@
Dense retrieval (Karpukhin et al., 2020) trains PLM-based encoders to condense queries and passages into low-dimensional dense vectors, then performs relevance search based on the Maximum Inner Product Search (MIPS) (Mussmann and Ermon, 2016) algorithm. In recent years, dense retrievers, such as BGE (Chen et al., 2024) and E5 (Wang et al., 2024b), have gained popularity for strong retrieval abilities across different tasks (Muennighoff et al., 2023; Thakur et al., 2021), languages (Zhang et al., 2023), and context granulaties (Williams et al., 2018; Nguyen et al., 2016), owing to diverse training data, retrieval-customized pre-training (Gao and Callan, 2021; Xiao et al., 2022; Ma et al., 2024), improved negative-mining (Xiong et al., 2021), and so on. Recent SOTA retrievers (Wang et al., 2024a; BehnamGhader et al., 2024; Lee et al., 2024) start to utilize LLMs (Jiang et al., 2023; Touvron et al., 2023) as backbone encoders. With larger parameters and pre-training data, LLM-based retrievers enable significantly powerful retrieval capacities and domain adaptation abilities. However, LLMs also incur heavy online workloads when serving as query encoders. Improving the serving efficiency of LLM-based retrievers is still left for exploration.
@@
Traditional statistical sparse retrieval algorithms (e.g., TF-IDF (Salton and Buckley, 1988), and BM25 (Robertson et al., 1994)) do not involve learnable language models. Instead, they directly analyze lexical frequencies (e.g., TF) and importance (e.g., IDF) in a large vocabulary space. These lexicals are sparsely distributed and require building inverted indexes (Luhn, 1957) for feasible similarity metric calculation, such as Apache Lucene (Bialecki et al., 2012) and Tantivy. In recent years, sparse retrieval has started to work with pre-trained LMs (PLMs). BGE (Chen et al., 2024), SPLADE (Formal et al., 2021), SparTerm (Bai et al., 2020), and SPARTA (Zhao et al., 2021) use customized XLM-RoBERTa (Conneau et al., 2020) or BERT (Devlin et al., 2019) as backbone encoders to learn term frequencies in an end-to-end manner and directly perform impact search based on the dot product of sparse vectors. Because of sparsity, these works also search with inverted indexes. Recent works (Zhuang et al., 2024) also try to incorporate LLMs with sparse retrieval. However, the online efficiency issue still exists. Learning dynamic analysis (Mackenzie et al., 2023) on learned sparse retreival shows that semantic information can be fully carried on the document side regardless of token semantics. Motivated by this finding, in this paper we hypothesize that if the semantic capacity is concentrated on the document side, then the query side can be made extremely lightweight with minimal performance degradation.
@@
Hybrid retrieval is a technique that combines retrieval scores (Wang et al., 2021) or ranks (Cormack et al., 2009) from multiple retrieval systems, such as dense (Karpukhin et al., 2020) / sparse (Robertson et al., 1994; Formal et al., 2021) retrieval, and single (Karpukhin et al., 2020) / multiple (Khattab and Zaharia, 2020; Chen et al., 2024) vector retrieval. It enables better retrieval abilities by interpolating results from multiple sources. Linear interpolation (Wang et al., 2021) of dense and sparse scores is used in our work.
@@ A.2.2 Improving inference efficiency
Previous works have tried to distill small query encoders from large encoders for better query inference efficiency. KALE (Campos et al., 2023) prunes the BERT-based query encoder and distills it during post-training. However, when the query encoder is pruned to one Transformers layer, its recall@20 on NQ (Kwiatkowski et al., 2019) decreases by 25-30%. (Wang and Lyu, 2023) proposes asymmetric dual-encoders by pruning BERT layers or using smaller BERT models. Its experiments show that a 2-layer BERT query encoder retains 92.5% of the full-sized performances via distillation and proper student initialization. USTAD (Kim et al., 2024) also proposes similar asymmetry and distillation strategies, which choose the 12-layer BERT as teachers and the 6-layer DistilBERT or 2-layer BERT-mini as students. Its experiments show that the distilled BERT-mini query encoder achieves 95-97% of the teacher’s performance because of better distillation.
However, these studies have primarily focused on small language models (SLMs), such as BERT (Devlin et al., 2019), and overlooked feasible approaches for improving online inference efficiency of LLM-based retrievers, where such optimization is even more critical. Furthermore, these existing methods typically retain Transformer layers in the query encoders, which limits their efficiency gains. In addition, most of these works rely on SLMs and are trained on relatively narrow datasets, such as MS-MARCO (Nguyen et al., 2016), thus unable to compete with LLM-based retrievers.
@@
Previous works have also been dedicated to reducing index sizes for smaller disk space usage. Matryoshka representation learning (MRL) (Kusupati et al., 2022) supports adjusting the dense vector dimensions by directly taking the top-k dimension of the original vector as shrink vectors. It aligns and trains the corresponding multi-dimensional vectors in the end-to-end contrastive learning. For example, if $k\in\{64,128,256\}$, query-passage dense vectors with $\{64,128,256\}$-dims could be obtained together after one encoder forward operation. For LM-based sparse retrieval, the similarity scores are directly affected by the impacts of corresponding lexicals (Formal et al., 2021). Thus, it naturally supports reducing the index sizes by retaining Top-k terms/lexicals. To explore the potential of controlling index sizes, our work also explores shrinking index sizes with dense MRL and sparse Top-k in the Section A.4, which also shows strong generalization capacities.
A.3 Case Study
Here, we report some hand-crafted ranking cases to show LightRetriever’s ability to understand complex queries rather than simple lexical matches, even if the query encoder is simplified and limited for inference efficiency.
Table 6: Case Study 1: Distinguish compositional semantics. Texts colored in red are important for semantic matching, while the blue ones represent unrelated words or distractions. $\checkmark$ means ranking correct, while $\times$ means rank fail.
Query Which camera captures fast-moving subjects with minimal motion blur?
Doc A (relevant) The X200 features a 1/8000s shutter and phase-detection AF, enabling crisp photos of sports and fast action with minimal motion blur.
Doc B The Y7 is praised for its wide-angle lens and superior low-light performance, often used for landscapes and night photography.
Results Full-sized Llama8b $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bDense $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bSparse $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bHybrid $\checkmark$ Doc A >> Doc B
Case Study 1: Distinguish compositional semantics.
This case tests the ability to search multiple attributes together, capturing fast-moving subjects with minimal motion blur. While both candidate documents discuss camera capabilities, only Doc A contains the correct composition of features. Doc B, although relevant to photography, emphasizes low-light and landscape use cases, which do not align with the compositional semantics. LightRetriever, despite using a simplified query encoder, successfully prioritizes Doc A across dense, sparse, and hybrid variants.
Table 7: Case Study 2: Negation.
Query Articles arguing against remote work for improving team collaboration.
Doc A (relevant) A recent study argues that remote work hinders spontaneous collaboration, noting dropoffs in serendipitous conversations and mentoring opportunities among junior staff.
Doc B Many experts advocate remote work for flexibility and increased individual productivity, highlighting asynchronous tools that maintain collaboration.
Results Full-sized Llama8b $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bDense $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bSparse $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bHybrid $\checkmark$ Doc A >> Doc B
Case Study 2: Negation.
This case tests whether models can correctly interpret affirmation/negation. The query explicitly seeks articles against remote work for improving collaboration. Although both candidate documents discuss remote work and collaboration, only Doc A aligns with the negative phase by describing how remote work hinders interactions. Doc B, on the contrary, advocates for remote work, which is semantically opposite to the query intent. Correctly distinguishing the affirmation/negation requires more than lexical overlap with remote work and collaboration. As shown in the results, both the full-sized Llama8b and all LightRetriever variants successfully rank Doc A higher, demonstrating that LightRetriever can preserve such negation discrimination ability.
Table 8: Case Study 3: Complex understanding.
Query How to revert the last Git commit without losing the changes?
Doc A (relevant) Use git reset –soft HEAD 1 to undo the last commit while keeping changes in the working tree and index; this retains your edited files for further edits.
Doc B Use git revert HEAD to create a new commit that undoes the changes introduced by the previous commit, preserving history and not altering the working tree.
Results Full-sized Llama8b $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bDense $\times$ Doc A << Doc B LightRetriever-Llama8bSparse $\checkmark$ Doc A >> Doc B LightRetriever-Llama8bHybrid $\times$ Doc A << Doc B
Case Study 3: Complex understanding.
This case tests the ability to understand complex technical instructions. The query asks how to revert the last Git commit without losing the changes, a subtle requirement that distinguishes between keeping edits in the working tree (Doc A) versus creating a new commit that reverts prior changes (Doc B). Despite both documents mentioning ways to undo commits, only Doc A satisfies the instructions. The full-sized Llama8b is able to complete such complex technical instructions. However, LightRetriever-Dense and Hybrid fail, likely because they focus on the strong lexical signal of revert but neglect the constraint without losing the changes. Interestingly, LightRetriever-Sparse ranks Doc A correctly, suggesting that token-level sparse matching may emphasize critical keywords such as keeping changes and working tree. This highlights a limitation of lightweight dense encoding: complex compositional understanding may fail, and success may depend on complementary sparse signals.
A.4 Additional results about controls of vector dimension, sparsity, and embedding serving size
Retrieval systems usually involve indexing millions of documents into search databases, which consumes large amounts of disk space, especially for dense vectors. For example, indexing 8.8M MS-MARCO passages (Nguyen et al., 2016) with the Llama-8b model consumes 134.92 GB for Faiss dense retrieval and 11.24 GB for Lucene sparse retrieval. Besides index sizes, LightRetriever’s dense part also serves as an Embedding matrix in RAM with shape [V, H] in Float16 precision. For the Llama-8b model, if the serving size is not properly controlled, such online embedding will consume 1.05 GB of RAM.
Thus, it’s essential to explore the possibility of controlling index sizes and serving sizes by shrinking the vector dimensions. 1) For dense retrieval, we incorporate Matryoshka Representation Learning (MRL) (Kusupati et al., 2022) with LightRetriever to support flexible controls of dense embedding dimensions. Given a dense vector $v^{den}$ with dimension $H$, MRL derives dimension-shrunk vectors by simply cutting out the Top-k dimensions $v^{den}[:k]$, then applies the same training loss as the original vectors. MRL trains a collection of Top-k dimensions together, enabling flexible choosing of a proper k. 2) For sparse retrieval, the sparse vector directly learns the term impact (i.e., lexical frequency) within a large sparse dimension. Thus, LightRetriever supports controlling sparsity by directly taking the Top-k terms, where the larger impact terms matter more.
@@ A.4 Additional results about controls of vector dimension, sparsity, and embedding serving size
Table 9: Performances, dense serving size, and index size of LightRetriever-Llama8b across different top-k dimensions (Dim). Note that dim=4096 is the untruncated setting.
Dimension Performance / nDCG@10 Serving Size (MB) Index Size (GB)
Top-k BEIR CMTEB-R Embedding Dense Sparse
128 49.3 55.2 32.83 4.22 3.03
256 51.6 57.5 65.67 8.43 5.42
512 53.0 58.9 131.33 16.87 8.89
1024 53.1 59.8 262.67 33.73 10.85
2048 53.3 60.4 525.34 67.46 11.22
4096 54.4 63.0 1050.67 134.92 11.24
Index size controls via tuning vector dimension and sparsity
We conduct experiments of controlling dense vector dimension and sparse vector sparsity with Top-k of $\{128,256,512,1024,2048,4096\}$ on Llama-3.1-8b. As presented in Figure 6 and Table 9, LightRetriever has affordable performance cost when truncating the Top-k. The lowest Top-k setting ($128$) has 5.1 and 7.8 nDCG@10 performance gaps on BeIR and CMTEB-R benchmarks compared to the highest Top-k ($4096$), but its disk usages with MS-MARCO are only 4.22 GB for dense trieval and 3.03 GB for sparse retrieval. This shows that our LightRetriever is capable of flexible index controls via simple vector shrinking techniques, with limited effect on retrieval performances.
Dense embedding serving size controls via dimension truncation
[…]
A.5 Additional ablation about the effect of auxiliary KL loss